plz helpCondensed balance sheet and income statement data for Pina Corporation are presented here. Pina Corporation Balance Sheets De
Answer #1
Particulars 2020 2019
Net Income (NI) 91000 66000
Net Revenue (R ) 705000 575000
Profit Margin = (NI / R) 12.91% 11.48%
Gross Profit (GP) 273000 218000
Gross Profit rate = (GP / R) 38.72% 37.91%
Opening Fixed Assets 370000 358000
Closing Fixed Assets 500000 370000
Average Fixed Assets (AFA) = [ (opening + closing) / 2 ] 435000 364000
Asset Turnover = (R / AFA) 1.62 1.58
Total Common stock, $10 per stock 327000 317000
Therefore, Total number of stock = (Total stock / 10) 32700 31700
Earnings per share EPS = (R / number of stock) 21.56 18.14
Market Price per Stock (MPS) 8.5 7.5
Price earning Ratio = (MPS / EPS) 0.39 0.41
Computation of Dividend Paid:
Opening Balance in Retained Earnings 127000 115000
Add: Current year earnings (i.e Net Revenue) 91000 66000
Less: Closing Balance in Retained Earnings 177000 127000
Therefore, Total Dividends paid 41000 54000
Total Number of stock 32700 31700
Dividend Per Stock (DPS) 1.25 1.70
Payout Ratio = (DPS / EPS) 5.82% 9.39%
Total Debt = Long Term Debt 152000 92000
Total Asset 743000 618000
Debt Asset Ratio = (Total Debt / Total Assets) 20.46% 14.89%